OpenStudy (anonymous):

How many solutions does the equation sin(5x) = 1/2 have on the interval (0, 2PI]

OpenStudy (anonymous):

Well, if you take a 30/60/90 triangle you get:  Therefore, when theta = 30 you get 1/2. 30 degrees = pi/6 Now then, you have to deal with the sign. They are positive in:  You would be correct in the first and second quadrants. pi/6 and 5pi/6 So 2 times.

OpenStudy (raden):

case I : sin5x=1/2 sin5x=sinpi/6 5x=(n*2pi+pi/6) x=(n*2pi/5+pi/30) just check for n=0,1,2,... case II : sin5x=1/2 sin5x=sin5pi/6 5x=(n*2pi+5pi/6) x=(n*2pi/5+pi/6 and again u must check for n=0,1,2,3,... which x's satisfied in interval (0,2pi]

OpenStudy (anonymous):

so..twice?

OpenStudy (anonymous):

RadEn What are you doing?

OpenStudy (raden):

yeps.. but sometimes there are duplicate number for x, so u just take once

OpenStudy (raden):

what do u mean @malical

OpenStudy (anonymous):

I don't know why you have N's.

OpenStudy (raden):

ohhh, it is a identity for trigono equation ^^

OpenStudy (anonymous):

sin5x=sinpi/6 5x=(n*2pi+pi/6) Why don't you just take the arcsin of each?

OpenStudy (anonymous):

er

OpenStudy (raden):

because the question to find "how many....." i think we didnt need arcsin of each, becasue we can get by manual to find for x's

OpenStudy (anonymous):

I'm stepping out of Spectrum's question and looking at your word directly.

OpenStudy (anonymous):

work* rather

OpenStudy (raden):

nopes.. @malical dont forget this one, one of identity for sin : sin x = sin(n*2pi + x)

OpenStudy (anonymous):

That is stupid. You are just adding 2pi which is a full circle. That's fine and all, but it isn't needed.

OpenStudy (raden):

have u tried it ?? @malical be carefull for this case : 5x=(n*2pi+pi/6) x=(n*2pi/5+pi/30) now, i give some examples for u. for n=0, it means x=pi/30 (satisfying for interval (0;2pi]) n=1, it means x=2pi/5+pi/30=13pi/30 (right too :p) . . . and so on, until u meet no satisfied of x in interval (0,2pi]

OpenStudy (anonymous):

What that means is that if you go around the unit circle (2pi) once you get the same answers. n2pi just means more unit circles. That is what the identity is. Example: sin[6pi] Is the same thing as sin[3(2pi)] And 2pi=0 So you get Sin[0].

OpenStudy (raden):

oke, i see that lets make a new question: sin2x=1/2 (x in interval [0,2pi] find the value of x ?

OpenStudy (anonymous):

pi/12 and 5pi/12

OpenStudy (raden):

what ur answer @malical

OpenStudy (raden):

oke, i will demonstrate it for u @malical sin2x=1/2 case I : sin2x=sin(pi/6) sin2x=sin(n*2pi + pi/6) 2x=n*2pi + pi/6 x=n*pi + pi/12 if n=0, then x=pi/12 if n=1, then x=pi + pi/12 = 13pi/12 (still satisfied) if n=2, then x=2pi + pi/12 = 25pi/12 (not satisfied) so on, for n=3,4,5,... not satisfied in interval (0,2pi] case II : sin2x=sin(5pi/6) sin2x=sin(n*2pi + 5pi/6) 2x=n*2pi + 5pi/6 x=n*pi + 5pi/12 if n=0, then x=5pi/12 if n=1, then x=pi+ 5pi/12 = 17pi/12 if n=2, then x=2pi+ 5pi/12=25pi/12 so on, for n=3,4,5... not satisied in interval (0,2pi] so, there are 4 solution

OpenStudy (raden):

the solution are : pi/12 , 13pi/12 , 5pi/12 , and 17pi/12

OpenStudy (raden):

well, @spectrum... i hope u can finish ur problem like i done

OpenStudy (raden):

i want to going to masque, know good luck .... Byeee

OpenStudy (raden):

*now

OpenStudy (anonymous):

It is 2. Because if it is x from 0 to 2pi that would work. But it is 2x from 0 to 2pi. So it is only 2.

OpenStudy (anonymous):

The other two would be -1/2

OpenStudy (anonymous):

sin2x if i plug x=13pi/12 gives sin(2*13pi/12) = sin390 degrees = sin30=1/2 not -1/2 looks raden is right

OpenStudy (anonymous):

yea, for x=17pi/12 sin(2*17pi/12) = sin510 degrees = sin150=1/2 too

OpenStudy (anonymous):

so how many are there ._.

OpenStudy (anonymous):

oh btw..it wasn't 4 ..it was 10 -.-

OpenStudy (raden):

4 was from my question, not ur problem @Spectrum :p

OpenStudy (anonymous):

note to self...do not skim for answer at 6:00 in the morning without sleep for awhile

OpenStudy (raden):

so, do u want its specific's answer

OpenStudy (anonymous):

i already got it wrong but you can help on my current question if you like it lolz

OpenStudy (raden):

haha... ok, what isit